if the ratio of two brothers' ages is 3 : 2
and their total age is 60 years old, what is the difference in their ages?

Answer by Maths68(1474)

Let
One Brother's age = x
Other Brother's age = y

The ratio of two brothers' ages is 3 : 2
x:y=3:2
x/y=3/2
x=3y/2..........(1)

Their total age is 60 years old
x+y=60
Put the value of x in above equation
3y/2+y=60
Multiply by 2 both sides
3y+2y=60*2
5y=120
5y/5=120/5
y=24
Put the value of y in (1)
x=3y/2
x=3(24)/2
x=3*12
x=36


One Brother's age = 36
Other Brother's age = 24
Difference between their ages = 36-24=12
Check
36:24=3:2
36/24=3/2
3/2=3/2